# Large-scale peculiar velocities in the universe

### v2

@@ 1 Introduction
- 6 Linear peculiar velocities
@@ 1.1 The enigma of bulk flows: what are they?
In standard cosmology, bulk flows are believed to be a natural consequence of structure formation. Tiny density perturbations in the primordial universe gradually grew through the physical mechanism of gravitational (or Jeans) instability, pulling matter towards overdense regions and gradually creating the large-scale structures we observe today, like the galaxy clusters and the superclusters. The enhanced gravitational pull from these structures generated the observed peculiar-velocity fields, driving the motion of the individual galaxies and creating the reported large-scale flows [5, 6]. The magnitude and direction of these flows provide a dynamic window to survey the distribution of both the luminous and the dark matter. However, despite a general consensus on the existence and the direction of these bulk motions, significant discrepancies still persist in their reported amplitudes and coherence scales. These inconsistencies, particularly the reports of extremely fast flows extending over unusually large distances, remain a source of contention and an active area of investigation within the cosmological community.
@@ 1.2 A persistent cosmological tension: bulk flows and the CDM model
While the existence of bulk flows is well-established, their precise amplitude and scale differ between the surveys, with some of them reporting bulk peculiar motions considerably faster and deeper than those predicted by the current cosmological model, namely by the CDM paradigm. Roughly speaking, the bulk flows reported by surveys confined within 100/h Mpc are generally in agreement with the CDM, while those going beyond the aforementioned threshold disagree. In fact, the increasing number of the latter surveys is becoming a significant and persistent problem for the standard cosmological model. The latter predicts that the amplitude of bulk flows should decrease with increasing scale and that the peculiar-velocity fields should converge to the rest-frame of the CMB beyond a depth of approximately 100/h Mpc. This expectation is largely based on Newtonian analyses, which predict a relatively slow growth-rate for linear peculiar velocities, scaling as [5, 6].
Numerous modern peculiar-velocity surveys have tested this prediction, yielding a mixed picture. As mentioned above, a number of studies find that the bulk flow amplitudes are in excellent agreement with the CDM paradigm, particularly on scales smaller than approximately 100/h Mpc (e.g. see [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23] for a representative list and also § 9.2 here). However, other surveys, often probing larger volumes, report bulk flows that are substantially faster and extend to greater depths than predicted (e.g. see [24, 25, 26, 27, 28, 29, 30, 31, 32] and § 9.3 here). Among them are measurements based on the recent Cosmic Flows 4 catalog, which revived the long-standing debate by reporting bulk flows in excess of the theoretical expectations on scales well beyond the 100/h Mpc mark [33, 34]. Even more dramatic are the reports of the controversial dark flows, which claim a coherent motion of galaxy clusters with velocities approaching 1000 km/sec on scales of several hundreds of megaparsecs, far exceeding the standard limits and potentially challenging the foundational principle of cosmic isotropy (see [35, 36, 37, 38, 39, 40] and also § 9.4 here).
Although not yet a definitive falsification of the CDM paradigm, these discrepancies highlight the limitations of our current understanding. The disagreement between the surveys may point to subtle biases in the observational techniques, to an underestimation of cosmic variance, or to the need for a more comprehensive theoretical framework that goes beyond the standard Newtonian approximation. Overall, the discrepancy underscores the importance of continued research into the bulk flows, positioning them as a critical test of the standard cosmological model and as a potential window into new developments.
@@ 1.3 Theoretical perspectives: Newtonian and relativistic views
Newtonian physics and general relativity are fundamentally different theories when it comes to gravity, because of the way they treat both the gravitational field and its sources. For example, in Newtonian theory the evolution of peculiar velocities is primarily driven by the gravitational acceleration generated by the emerging mass overdensities. This is the foundation of the gravitational instability paradigm, which is central to the standard CDM model. The equations of motion are based on a linear approximation of the gravitational potential and predict a relatively slow growth-rate for peculiar velocities after equipartition and during the matter-dominated era. This growth is generally understood to scale with time as , a result that has been consistently recovered in both the traditional Newtonian studies [5, 6] and in their quasi-Newtonian counterparts [41, 42] (see also § 6.1 and § 6.2 respectively). The aforementioned moderate growth-rate serves as the theoretical benchmark against which the observed amplitudes and scales of bulk flows are compared. Therefore, when a survey reports bulk velocities significantly in excess of these predictions, it signals a potential tension with the standard model.
@@
The relativistic approach, however, introduces a crucial refinement. According to general relativity, gravity is sourced not only by mass density, but also by matter fluxes (i.e. by matter in motion). This adds a non-Newtonian flux-related term to the energy-momentum tensor, which acts as an additional source of gravity and couples back into the equations governing the evolution of linear peculiar velocities. As demonstrated in a series of covariant relativistic studies (see [43, 44, 45, 46, 47] and also § 6.3 here), including the gravitational input of the peculiar flux leads to a considerably faster linear growth-rate for the peculiar-velocity field. More specifically, on a flat and matter-dominated Friedmann-Robertson-Walker (FRW) background universe, the minimum linear growth-rate of the relativistic analysis is , which is substantially stronger than the Newtonian prediction. This result offers a natural general-relativistic mechanism to explain the existence of high-amplitude bulk flows, potentially bridging the gap between theory and observation without requiring any new physics beyond the standard model. In this context, the tension between theory and observation is not a sign of a flawed model, but rather a reflection of the limitations of the Newtonian approximations employed to predict the growth of peculiar velocities. The debate thus shifts, from questioning the standard model itself, to scrutinizing the validity of the approximations used in its application to cosmic dynamics.
@@ 1.4 Methodological challenges and observational biases
A primary challenge lies in the intrinsic scatter of redshift-independent distance indicators, such as the Tully-Fisher (TF) relation and the Fundamental Plane (FP). While these relations provide robust distance estimates, they possess a finite intrinsic scatter that can lead to significant distance errors, particularly for individual galaxies. This uncertainty translates directly into the derived peculiar velocities, where a small fractional error in distance can result in a large absolute velocity error, especially at greater distances [48].
Furthermore, the process of galaxy selection and measurement introduces significant biases that must be rigorously addressed. The most prominent of these is the Malmquist bias, which arises in flux-limited surveys. There, intrinsically fainter galaxies are observed only if they are relatively nearby, while brighter galaxies can be seen at much greater distances. This leads to a systematic overestimation of distances for a given observed redshift. A related effect, known as the inhomogeneous Malmquist bias, occurs when peculiar motions themselves conspire with the density field. For example, galaxies in an overdense region are typically falling inward, making them appear closer than they are, which can create a spurious, large-scale infall signal in the reconstructed velocity field [49, 50].
@@ 1.5 Broader implications: from structure formation to cosmological tensions
Beyond this, the effects of relative motion can introduce apparent (Doppler like) anisotropies in the cosmological observations. In a “tilted universe”, where observers have a peculiar velocity relative to the cosmic rest-frame, their measurements will appear to be direction dependent. This dependence can manifest itself in a host of ways. For example, as a dipolar anisotropy in the number counts of distant astrophysical sources, or in the measurements of cosmological parameters. In fact, such dipoles have been reported in the number counts of radio galaxies and quasars [61, 62, 63, 64, 65, 66, 67, 68], in the value of the Hubble constant [57, 58, 59], in the measurements of the universal acceleration [69, 70, 71, 72] and in the distribution of [73]. Moreover, the reported dipoles seem consistent with our local bulk flow. Although the cosmic-dipole question has not been settled yet (e.g. see [74, 75, 76, 77] for a “sample” of related discussions), the findings of [71, 72, 73], combined with the theoretical predictions of [78, 79, 80, 81], raise the profound question whether a portion (at least) of the observed cosmic acceleration could be a local kinematic illusion rather than a global cosmological event driven by dark energy? After all, the history of astronomy is replete with examples of relative motions leading to profound misinterpretations, from planetary retrograde motion to the initial understanding of the CMB dipole, underscoring the need for continued vigilance in separating local kinematic effects from genuine cosmic phenomena. Having said that, the same aforementioned findings could also be seen as a challenge to the “Cosmological Principle” and to the very foundations of modern cosmology [66, 82].
@@ 1.6 Objective and scope of the review
Structure formation and the nonlinear regime: The role of peculiar velocities during structure formation is still rather poorly understood, since the available studies are few and sparse and all Newtonian/quasi-Newtonian in nature. According to the latter, the linear evolution of structure proceeds as in the absence of peculiar-velocity perturbations. Also, employing the quasi-Newtonian approach to revisit the Zel’dovich approximation [56], which applies to the mildly nonlinear phase of structure formation, simply recovered the standard Newtonian “pancake” attractor (see [86], as well as § 8.4 here). The Newtonian picture was also fully recovered in the relativistic treatment of the “Meszaros stagnation effect” [87], though only after the gravitational contribution of the peculiar flux was switched off by moving the study to the Landau-Lifshitz frame (see [88, 89] and also § 8.3 here).
@@ 2.1.2 Initial systematic surveys
The evolution of distance indicators proved crucial for understanding large-scale flows. The Fundamental Plane relation emerged as a powerful tool, establishing that elliptical galaxies occupy a well-defined plane in the space of their physical parameters. This method achieved distance uncertainties of around 17% for individual galaxies, representing a significant improvement over earlier techniques. Major surveys like EFAR and SMAC employed this relation to map flows extending to 15,000 km/sec, revealing complex patterns of large-scale motion.
For spiral galaxies, the Tully-Fisher relation provided a complementary approach, correlating galaxy luminosity with rotation velocity. Modern implementations, particularly in the infrared, achieved typical distance uncertainties of 15-20%. This method proved especially valuable for mapping flows in less dense regions and through the Zone of Avoidance, where traditional optical techniques struggle.
@@ 2.2.3 The Cosmic Flows program progression (CF1 to CF4)
The latest iteration, Cosmic Flows-4 (CF4) [85], has expanded to include distances for 55,877 galaxies gathered into 38,065 groups. The most important addition is a much-extended T-F sample of 10,000 galaxies drawing particularly on kinematic information from ALFALFA, the Arecibo Legacy Fast ALFA survey of the high galactic latitude sky in the declination range 0°-38°. Photometry is provided by the Sloan Digital Sky Survey (SDSS) and the Wide-field Infrared Explorer (WISE). This component substantially redresses the imbalance favoring the southern sky of the previous catalog.
@@
Each release has employed eight methodologies: Tully-Fisher, Fundamental Plane, Surface Brightness Fluctuations (SBF), Type Ia supernovae, Type II supernovae, Tip of the Red Giant Branch (TRGB), Cepheid period-luminosity relation, and nuclear masers. The assembly of galaxies into groups is an important feature of the study in facilitating overlaps between methodologies. Merging between multiple contributions within a methodology and between methodologies is carried out with Bayesian Markov chain Monte Carlo procedures. The final assembly of distances is compatible with a value of the Hubble constant of km/secMpc with a statistical error of ±0.8 km/secMpc, but with a large potential systematic error of 3 km/secMpc as well.
@@ 2.3.1 Consensus on the bulk-flow direction
The robustness of this directional consensus is demonstrated by strong agreement across diverse measurement techniques. The 2MTF survey found a bulk flow pointing toward [14], while the 6dFGSv analysis indicated a nearly identical direction of [17]. The recent CF4 results further strengthen this consensus, showing alignment with [85]. Even the latest Pantheon+ SNIa analysis, using entirely different methodology, finds a consistent direction of [135].
@@ 2.3.4 Systematic effects and biases
The measurement of bulk flows is subject to numerous systematic effects and biases that can significantly impact the interpretation of observational data. Understanding and correctly accounting for these systematic effects has become crucial for obtaining reliable results and reconciling apparently conflicting measurements.
@@
Malmquist bias represents one of the most significant systematic effects in peculiar velocity surveys. This bias manifests itself in two distinct forms: homogeneous and inhomogeneous Malmquist bias [48]. The homogeneous form arises from the coupling between distance errors and the galaxy number density function, while the inhomogeneous variant stems from local density variations along the line of sight. Recent work by Davis et al [49] has shown that failure to properly account for these effects can lead to systematic errors in the bulk-flow measurements of up to 150 km/sec.
@@
Selection effects present another major challenge, particularly in magnitude-limited surveys. Hudson and Turnbull [50] demonstrated that selection criteria based on apparent magnitude can introduce artificial patterns in the measured velocity field, especially at the survey boundaries. These effects become more pronounced at larger distances, where selection functions typically become more stringent.
@@
The zone of avoidance, caused due to obscuration by the Galactic plane, introduces significant complications in the bulk-flow measurements. While various interpolation schemes have been proposed [140], the missing data in this region can affect both the magnitude and direction of measured bulk flows. Studies by Erdoğdu et al [51] suggest that incomplete sky coverage can bias the bulk-flow measurements by up to 10-15% in magnitude and 5-10 degrees in direction.
@@
Distance indicator calibration represents another crucial source of systematic uncertainty. Each distance-indicator method carries its own systematic effects: Tully-Fisher relations suffer from rotation-curve asymmetries and inclination effects; Fundamental Plane measurements are affected by population gradients and environmental dependencies; and Surface Brightness Fluctuation methods show sensitivity to stellar population variations. These effects can introduce correlated errors in peculiar velocity measurements across entire surveys.
@@
Non-linear effects in the velocity field pose additional challenges. While linear theory provides a good approximation on large scales, smaller scale non-linear motions can contaminate the bulk-flow measurements. Kaiser and Hudson [141] developed a framework for treating these non-linear effects, showing that they can bias the bulk-flow measurements especially in dense environments and at small scales.
@@
Proper motion measurement and correction present another layer of complexity. Recent work by Carrick et al [15] has shown that coherent stellar streaming motions can contaminate galaxy peculiar-velocity measurements, particularly for nearby galaxies where proper motions are significant. This effect requires careful consideration in modern surveys attempting to map local flow fields.
@@
The treatment of galaxy groups and clusters introduces additional systematic considerations. While grouping algorithms help reduce noise in peculiar-velocity measurements, they can also introduce biases if the group membership is incorrectly assigned, or if the group dynamics are not properly modeled. Recent work has focused on developing more sophisticated grouping algorithms that account for these effects while preserving the underlying flow signal.

### v3

@@ 1 Introduction
6 Linear peculiar velocities
- 6.1 Newtonian analysis
- 6.2 Quasi-Newtonian analysis
6.3 Relativistic analysis
- 6.3.1 Two alternative perspectives
- 6.3.2 The key role of the peculiar flux
- 6.3.3 The relativistic 4-acceleration
- 6.3.4 Peculiar velocities in FLRW universes with
- 6.3.5 The role of the peculiar 4-acceleration
- 6.3.6 Peculiar velocities in FLRW universes with
- 6.3.7 Peculiar velocities before recombination
- 6.4 Peculiar velocities and inflation
- 6.5 Peculiar velocities and dark-energy
- 6.6 Cosmologies with a generic “tilt”
@@ 1.1 The enigma of bulk flows: what are they?
In standard cosmology, bulk flows are believed to be a natural consequence of structure formation. Tiny density perturbations in the primordial universe gradually grew through the physical mechanism of gravitational (or Jeans) instability, pulling matter towards overdense regions and gradually creating the large-scale structures we observe today, like the galaxy clusters and the superclusters. The enhanced gravitational pull from these structures generated the observed peculiar-velocity fields, driving the motion of the individual galaxies and creating the reported large-scale flows [5, 6]. The magnitude and direction of these flows provide a dynamic window to survey the distribution of both the luminous and the dark matter. However, despite a general consensus on the existence and the direction of these bulk motions, significant discrepancies still persist in their reported amplitudes and coherence scales. These inconsistencies, particularly the reports of extremely fast flows extending over unusually large distances, remain a source of contention and an active area of investigation within the cosmological community. A detailed discussion of these discrepancies and the tension they create with the standard model is presented in § 9.3 and § 9.4.
@@ 1.2 A persistent cosmological tension: bulk flows and the CDM model
While the existence of bulk flows is well-established, their precise amplitude and scale remain a subject of debate, with reports of peculiar motions often deviating from the standard CDM paradigm. Theoretically, based largely on Newtonian analyses where linear peculiar velocities grow slowly as [5, 6], the standard model predicts that bulk flow amplitudes should decrease with scale, converging to the CMB rest-frame beyond approximately 100/h Mpc. Testing this prediction has yielded a mixed picture that roughly splits according to the depth of the survey. Numerous studies probing scales within the 100/h Mpc threshold find amplitudes in excellent agreement with CDM expectations (e.g. see [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23] and § 9.2 here). Conversely, an increasing number of surveys probing larger volumes report flows that are substantially faster and deeper than predicted, posing a significant and persistent challenge to the standard model (e.g. see [24, 25, 26, 27, 28, 29, 30, 31, 32] and § 9.3 here). Notable among these are the recent Cosmic Flows 4 results, which revived the controversy by finding excessive bulk flows well beyond 100/h Mpc [33, 34], and the even more dramatic reports of dark flows, which claim coherent cluster motions approaching 1000 km/sec on scales of several hundreds of megaparsecs, potentially challenging the foundational principle of cosmic isotropy (see [35, 36, 37, 38, 39, 40] and also § 9.4 here).
Although not yet a definitive falsification of the CDM paradigm, these discrepancies highlight the limitations of our current understanding. The disagreement between the surveys may point to subtle biases in the observational techniques, to an underestimation of cosmic variance, or to the need for a more comprehensive theoretical framework that goes beyond the standard Newtonian approximation. Overall, the discrepancy underscores the importance of continued research into the bulk flows, positioning them as a critical test of the standard cosmological model and as a potential window into new developments. The disagreement between the surveys may point to subtle biases in the observational techniques (discussed in § 3), to an underestimation of cosmic variance, or to the need for a more comprehensive theoretical framework that goes beyond the standard Newtonian approximation (see § 6). Furthermore, systematic errors arising from instrumentation, such as calibration drifts, can mimic large-scale flows if not properly characterized (e.g. see [41] for Planck systematics).
@@ 1.3 Theoretical perspectives: Newtonian and relativistic views
Newtonian physics and general relativity are fundamentally different theories when it comes to gravity, because of the way they treat both the gravitational field and its sources. For example, in Newtonian theory the evolution of peculiar velocities is primarily driven by the gravitational acceleration generated by the emerging mass overdensities. This is the foundation of the gravitational instability paradigm, which is central to the standard CDM model. The evolution of peculiar velocities is typically derived using linearized Newtonian perturbation theory, where the gravitational potential is determined by the Poisson equation sourced by density fluctuations. This approach predicts a relatively slow growth-rate for peculiar velocities after equipartition and during the matter-dominated era. This growth is generally understood to scale with time as , a result that has been consistently recovered in both the traditional Newtonian studies [5, 6] and in their quasi-Newtonian counterparts. The latter are irrotational perturbed spacetimes with vanishing shear and magnetic Weyl tensor, that is with no gravitational waves, while their 4-acceleration is given by the gradient of an arbitrary (Newtonian-like) scalar potential. All these constraints hold at the linear perturbative level (see [42, 43] and also § 6.1 and § 6.2 in this review). The aforementioned moderate growth-rate serves as the theoretical benchmark against which the observed amplitudes and scales of bulk flows are compared. Therefore, when a survey reports bulk velocities significantly in excess of these predictions, it signals a potential tension with the standard model.
@@
The relativistic approach, however, introduces a crucial refinement. According to general relativity, gravity is sourced not only by mass density, but also by matter fluxes (i.e. by matter in motion). Unlike Newtonian physics, where momentum fluxes also exist but do not source the Poisson equation, in relativity the fluxes add to the energy-momentum tensor and act as an additional source of gravity, which then couples back to the equations governing the evolution of linear peculiar velocities.111For discussions on post-Newtonian studies and non-linear extensions, we refer the reader to [44, 45, 46]. As demonstrated in a series of covariant relativistic studies (see [47, 48, 49, 50, 51] and also § 6.3 here), including the gravitational input of the peculiar flux leads to a considerably faster linear growth-rate for the peculiar-velocity field. More specifically, on a flat and matter-dominated Friedmann-Lemaître-Robertson-Walker (FLRW) background universe, the minimum linear growth-rate of the relativistic analysis is , which is substantially stronger than the Newtonian prediction. This result offers a natural general-relativistic mechanism to explain the existence of high-amplitude bulk flows, potentially bridging the gap between theory and observation without requiring any new physics beyond the standard model. In this context, the tension between theory and observation is not a sign of a flawed model, but rather a reflection of the limitations of the Newtonian approximations employed to predict the growth of peculiar velocities. The debate thus shifts, from questioning the standard model itself, to scrutinizing the validity of the approximations used in its application to cosmic dynamics.
@@ 1.4 Methodological challenges and observational biases
A primary challenge lies in the intrinsic scatter of redshift-independent distance indicators, such as the Tully-Fisher (TF) relation [52] and the Fundamental Plane (FP) [53]. While these relations provide robust distance estimates, they possess a finite intrinsic scatter that can lead to significant distance errors, particularly for individual galaxies. This uncertainty translates directly into the derived peculiar velocities, where a small fractional error in distance can result in a large absolute velocity error, especially at greater distances [54].
Furthermore, the process of galaxy selection and measurement introduces significant biases that must be rigorously addressed. The most prominent of these is the Malmquist bias [55], which arises in flux-limited surveys. There, intrinsically fainter galaxies are observed only if they are relatively nearby, while brighter galaxies can be seen at much greater distances. This leads to a systematic overestimation of distances for a given observed redshift. A related effect, known as the inhomogeneous Malmquist bias, occurs when peculiar motions themselves conspire with the density field to modulate the number of visible sources [54]. For example, galaxies in an overdense region are typically falling inward, making them appear closer than they are, which can create a spurious, large-scale infall signal in the reconstructed velocity field [56, 57].
@@ 1.5 Broader implications: from structure formation to cosmological tensions
Beyond this, the effects of relative motion can introduce apparent (Doppler like) anisotropies in the cosmological observations. In a “tilted universe”, where observers have a peculiar velocity relative to the cosmic rest-frame, their measurements will appear to be direction dependent. This dependence can manifest itself in a host of ways. For example, as a dipolar anisotropy in the number counts of distant astrophysical sources, or in the measurements of cosmological parameters. In fact, such dipoles have been reported in the number counts of radio galaxies and quasars [68, 69, 70, 71, 72, 73, 74, 75] (see also [76, 77, 78, 79] for discussions on the kinematic origin of the dipole), in the value of the Hubble constant [64, 65, 66], in the measurements of the universal acceleration [80, 81, 82, 83] and in the distribution of [84]. Moreover, the reported dipoles seem consistent with our local bulk flow. Although the cosmic-dipole question has not been settled yet (e.g. see [85, 86, 87, 88] for a “sample” of related discussions), the findings of [82, 83, 84], combined with the theoretical predictions of [89, 90, 91, 92], raise the profound question whether a portion (at least) of the observed cosmic acceleration could be a local kinematic illusion rather than a global cosmological event driven by dark energy? After all, the history of astronomy is replete with examples of relative motions leading to profound misinterpretations, from planetary retrograde motion to the initial understanding of the CMB dipole, underscoring the need for continued vigilance in separating local kinematic effects from genuine cosmic phenomena. Having said that, the same aforementioned findings could also be seen as a challenge to the “Cosmological Principle” and to the very foundations of modern cosmology [73, 93].
@@ 1.6 Objective and scope of the review
Structure formation and the nonlinear regime: The role of peculiar velocities during structure formation is still rather poorly understood, since the available studies are few and sparse and all Newtonian/quasi-Newtonian in nature. According to the latter, the linear evolution of structure proceeds as in the absence of peculiar-velocity perturbations. Also, employing the quasi-Newtonian approach to revisit the Zel’dovich approximation [63], which applies to the mildly nonlinear phase of structure formation, simply recovered the standard Newtonian “pancake” attractor (see [97], as well as § 8.4 here). The Newtonian picture was also fully recovered in the relativistic treatment of the “Meszaros stagnation effect” [98], though only after the gravitational contribution of the peculiar flux was switched off by moving the study to the Landau-Lifshitz frame. By default, the latter coordinate system has zero energy flux but nonzero particle flux (see [99, 100] and also § 8.3 here).
@@ 2.1.2 Initial systematic surveys
The evolution of distance indicators proved crucial for understanding large-scale flows. The Fundamental Plane relation emerged as a powerful tool, establishing that elliptical galaxies occupy a well-defined plane in the space of their physical parameters. This method achieved distance uncertainties of around 17% for individual galaxies, representing a significant improvement over earlier techniques. Major surveys, such as EFAR [105] and SMAC [24], employed this relation to map flows extending to 15,000 km/sec, revealing complex patterns of large-scale motion.
For spiral galaxies, the Tully-Fisher relation provided a complementary approach, correlating galaxy luminosity with rotation velocity. Modern implementations, particularly in the infrared, achieved typical distance uncertainties of 15-20%. This method proved especially valuable for mapping flows in less dense regions and through the “Zone of Avoidance”, where traditional optical techniques struggle due to high extinction [106].
@@ 2.2.3 The Cosmic Flows program progression (CF1 to CF4)
The latest iteration, Cosmic Flows-4 (CF4) [96], has expanded to include distances for 55,877 galaxies gathered into 38,065 groups. The most important addition is a much-extended T-F sample of 10,000 galaxies drawing particularly on kinematic information from ALFALFA, the Arecibo Legacy Fast ALFA survey of the high galactic latitude sky in the declination range . Photometry is provided by the Sloan Digital Sky Survey (SDSS) and the Wide-field Infrared Explorer (WISE). This component substantially redresses the imbalance favoring the southern sky of the previous catalog.
@@
Each release has employed eight methodologies: Tully-Fisher, Fundamental Plane, Surface Brightness Fluctuations (SBF), Type Ia supernovae, Type II supernovae, Tip of the Red Giant Branch (TRGB), Cepheid period-luminosity relation, and nuclear masers. The assembly of galaxies into groups is an important feature of the study in facilitating overlaps between methodologies. Merging between multiple contributions within a methodology and between methodologies is carried out with Bayesian Markov chain Monte Carlo procedures. The final assembly of distances is compatible with a value of the Hubble constant of km/secMpc with a statistical error of km/secMpc, but with a large potential systematic error of 3 km/secMpc as well.
@@ 2.3.1 Consensus on the bulk-flow direction
The robustness of this directional consensus is demonstrated by strong agreement across diverse measurement techniques. The 2MTF survey found a bulk flow pointing toward [14], while the 6dFGSv analysis indicated a nearly identical direction of [17]. The recent CF4 results further strengthen this consensus, showing alignment with [96]. Even the latest Pantheon+ SNIa analysis, using entirely different methodology, finds a consistent direction of [146]. These results are further supported by extensive photometric and spectroscopic data from SDSS [147] and WISE [148], which have provided the large-scale structure maps necessary to interpret these flows.
@@ 2.3.4 Systematic effects and biases
The measurement of bulk flows is subject to numerous systematic effects and biases that can significantly impact the interpretation of observational data. Understanding and correctly accounting for these systematic effects has become crucial for obtaining reliable results and reconciling apparently conflicting measurements. Key systematic challenges include:
@@
Malmquist bias: This represents one of the most significant systematic effects in peculiar velocity surveys. This bias manifests itself in two distinct forms: homogeneous and inhomogeneous Malmquist bias [54]. The homogeneous form arises from the coupling between distance errors and the galaxy number density function, while the inhomogeneous variant stems from local density variations along the line of sight. Recent work by Davis et al [56] has shown that failure to properly account for these effects can lead to systematic errors in the bulk-flow measurements of up to 150 km/sec.
@@
Selection effects: These present another major challenge, particularly in magnitude-limited surveys. Hudson and Turnbull [57] demonstrated that selection criteria based on apparent magnitude can introduce artificial patterns in the measured velocity field, especially at the survey boundaries. These effects become more pronounced at larger distances, where selection functions typically become more stringent.
@@
The Zone of Avoidance (ZoA): Caused due to obscuration by the Galactic plane, the ZoA introduces significant complications in the bulk-flow measurements. While various interpolation schemes have been proposed [106], the missing data in this region can affect both the magnitude and direction of measured bulk flows. Studies by Erdoğdu et al [58] suggest that incomplete sky coverage can bias the bulk-flow measurements by up to 10-15% in magnitude and 5-10 degrees in direction.
@@
Distance indicator calibration: This represents another crucial source of systematic uncertainty. Each distance-indicator method carries its own systematic effects: Tully-Fisher relations suffer from rotation-curve asymmetries and inclination effects; Fundamental Plane measurements are affected by population gradients and environmental dependencies; and Surface Brightness Fluctuation methods show sensitivity to stellar population variations. These effects can introduce correlated errors in peculiar velocity measurements across entire surveys.
@@
Non-linear effects: In the velocity field, these pose additional challenges. While linear theory provides a good approximation on large scales, smaller scale non-linear motions can contaminate the bulk-flow measurements. For scales smaller than the shell-crossing scale, the velocity distribution becomes highly non-trivial due to phase mixing and multi-streaming regions. While standard linear theory breaks down here, requiring N-body simulations or kinetic theory approaches (e.g. see [44]), our focus remains on the large-scale bulk flows where linear and quasi-linear approximations are traditionally applied. Kaiser and Hudson [153] developed a framework for treating these non-linear effects, showing that they can bias the bulk-flow measurements especially in dense environments and at small scales.
@@
Proper motion measurement: Correction of proper motions presents another layer of complexity. Recent work by Carrick et al [15] has shown that coherent stellar streaming motions can contaminate galaxy peculiar-velocity measurements, particularly for nearby galaxies where proper motions are significant. This effect requires careful consideration in modern surveys attempting to map local flow fields.
@@
Galaxy grouping algorithms: The treatment of galaxy groups and clusters introduces additional systematic considerations. While grouping algorithms help reduce noise in peculiar-velocity measurements, they can also introduce biases if the group membership is incorrectly assigned, or if the group dynamics are not properly modeled. Recent work has focused on developing more sophisticated grouping algorithms that account for these effects while preserving the underlying flow signal.